The hailstorm (with about an inch diameter sized hailstones) yesterday afternoon curbed our hope for a clear night, so we used the afternoon for a new look at the noise characteristics, and we can confirm a similar read noise level as we did in the lab in Tucson. So far so good.
Todd took a first shot at the photometric zeropoints in the four ODI filters, and they are within 0.1 mag of the prediction we made a few years ago. As time permits, we will update the ODI exposure time calculator.
Surprisingly, the sky cleared up before 10pm, giving us the opportunity to spend more time to tune telescope guiding. A few hours of debugging and optimizing later, we ended up with a workable configuration of guide parameters that allows more stable guiding. Getting the guider fully tuned to a critically damped state will take some time, though.
In one instance we recorded a single bright star at a rate of >30Hz. While the telescope guider filters that signal to a slower rate, it also calculates the power spectrum of the image motion (the panels with the green histogram in the picture; ignore the units on the Frequency axis - they are meaningless at this time). Most of the time that power spectrum is boring, but when some wind blew into the dome, a clear spike showed up in the power spectrum, indicating that the telscope mount might have gone into its 8 Hz resonance. With still some bandwidth left in the system, this measurements is encouraging for the prospects of the coherently corrected OT modes with pODI, where this resonant image motion could be compensated in the detectors.